OpenStudy (anonymous):

A 10-kilogram box is at static equilibrium, and the downward pull of gravity acting on the box is 98 newtons. What is the minimum force that would be required to just pick up the box?

OpenStudy (anonymous):

F = m*a = 10*98 = 980